OpenStudy (anonymous):

use this form h(t) = -a (t-10)^2+100a to find the time after being fired out the cannon that the ball obtain maximum height. the maximum height of the cannon ball is 20m (above the end of the barrel which is 4.2 m high)

sam (.sam.):

What's the problem? If you find h'(t) , at max height h'(t) is zero

OpenStudy (anonymous):

which height must you use to get the time when the ball has reached the max height

sam (.sam.):



sam (.sam.):

Use 20m

OpenStudy (anonymous):

20=-a(t-10)^2 +100a is it correct the first step?

OpenStudy (anonymous):

or is it 20t

sam (.sam.):

Differentiate it first, because h'(t) is velocity, at max height, velocity is zero

OpenStudy (anonymous):

answer is 20s?

sam (.sam.):

Nope, \[h(t)=-a(t-10)^2+100a \\ \\ h'(t)=-2a(t-10)=0 \\ \\ t=10s\]

OpenStudy (anonymous):

why -2a?

sam (.sam.):

You're differentiating -a(t-10)^2, using power rule

OpenStudy (anonymous):

ok thanks a lot for all the trouble

sam (.sam.):

Ok so you got the answer?

OpenStudy (anonymous):

yes thanks 10?

sam (.sam.):

yes